OpenStudy (anonymous):

(3,-6); 7x+2y=9 write the equation of the line containing the given point and the perpendicular to the given line.

OpenStudy (aripotta):

what's the slope of 7x + 2y = 9?

OpenStudy (anonymous):

well it wants the answer in point slope form.

OpenStudy (aripotta):

ok, but you have to find the slope of the first line first before you can do anything else

OpenStudy (anonymous):

thats all my paper has im stuck in this question. unless the slope would be 3,-6 thats my guess.

OpenStudy (anonymous):

wait unless its y+6= 2/7 (x-3)

OpenStudy (aripotta):

that's right..

OpenStudy (aripotta):

that's your answer lol. good job

OpenStudy (anonymous):

na no way really.

OpenStudy (aripotta):

serious, bro. idk how you got it so fast, but BAM. there ye go

OpenStudy (anonymous):

lol but i dont even know how to start with the first step.

OpenStudy (anonymous):

well i wanna know to get the answer.

OpenStudy (anonymous):

would be like y=mx+b type thing.

OpenStudy (aripotta):

ok, so you change 7x + 2y = 9 into slope-intercept form. 7x + 2y = 9 2y = -7x + 9 y = -7/2x + 9/2

OpenStudy (aripotta):

and the slope in y = -7/2x + 9/2 is -7/2

OpenStudy (aripotta):

so if you want the slope of the line that's perpendicular to this, you find the negative reciprocal of -7/2, which is 2/7

OpenStudy (aripotta):

point-slope form is: y - y1 = m(x - x1) our y1 is -6, x1 is 3, and m is 2/7. so... y + 6 = 2/7(x - 3)

OpenStudy (anonymous):

wow dude the answer was in my notes but i was like no way thats why i wanted to know the steps to the answer thx.

OpenStudy (aripotta):

no problem :) i hope that made sense